Problem 9-11
Cost of Equity

Radon Homes's current EPS is $7.80. It was $4.43 5 years ago. The company pays out 50% of its earnings as dividends, and the stock sells for $30.

  1. Calculate the historical growth rate in earnings. (Hint: This is a 5-year growth period.) Round your answer to two decimal places.
         %
  2. Calculate the next expected dividend per share, D1 (Hint: D0 = 0.50($7.80) = $3.90). Assume that the past growth rate will continue. Round your answer to the nearest cent.
    $  
  3. What is Radon's cost of equity, rs? Round your answer to two decimal places     %
  4. Please show all formulas and steps to arrive at the solution.
Answer #1

a) growth in earnings, g = (D5 / D0)^(1/5) - 1 = (7.8 / 4.43)^(1/5) - 1 = 11.98%

b) D1 = D0 x (1 + g) = 3.90 x (1 + 11.98%) = $4.37

c) Cost of equity, rs = D1 / P + g = 4.37 / 30 + 11.98% = 26.54%